OpenStudy (anonymous):

explain how the graph of each given function is a transformation of the graph y=x^2 A.y=x^2-5

OpenStudy (anonymous):

Shift it down 5.

OpenStudy (anonymous):

So you can see the difference: http://www.wolframalpha.com/input/?i=plot+y%3Dx^2%2Cy%3Dx^2-5

OpenStudy (anonymous):

B.y=-2x^2 c.Y=(x-5)^2 d.y=(-3x)^2

OpenStudy (anonymous):

B. the - rotates it around the x axis and the 2 stretches it vertically C. the -5 INSIDE the ()'s moves it to the left 5 D. the - INSIDE rotates it around the y axis and the 3 stretches it (though once you square it, its a nine)